OpenStudy (anonymous):

write each fraction in simplest form if the fraction is already in simpliest form , write simplified 1.) 24/36

OpenStudy (anonymous):

2/3

OpenStudy (anonymous):

What about this one 30/42 ?

OpenStudy (unklerhaukus):

divide denominator and numerator by six

OpenStudy (anonymous):

i dont know how to do that :9

OpenStudy (unklerhaukus):

\[30 \div 6= 5\] \[42 \div 6= 7\] so\[\frac{30}{42}=\frac{1 \over 6}{1 \over 6}\frac{30}{42}=\frac{5}{7}\]

OpenStudy (anonymous):

 28x